Fermi National Laboratory
 

Fermilab Shutdown Report for August 21, 2009 - August 28, 2009

 

Fermilab Shutdown Report
8/21/09 – 8/28/09

LINAC

Shutdown Week 11 – On Schedule

  • Preaccelerator
    • H- Source is operational @ 55.5 mA
    • Repaired HQ6HQ10 supply (output drifting)
    • Worked with FESS to repair the HVAC unit for the Preaccelerator
    • I- Source failed (05/25) due to excessive sparks
      • Hipot showed that the next to last gap would only hold 80 kV as opposed to 125 kV
    • Extensive work was done during the shutdown to clean the ceramics
  • I - Source Status:
    • Started slowly HV conditioning on 08/10.
    • Ran @ 750 keV for almost 15 days. However
      • Excessive sparks occurred this weekend
      • Latest hipot of the same gap revealed that is not holding above 55 kV
    • Current Plan:
      • Let up the I- column
      • Remove the electrodes
      • Evaluate the situation
    • Vulnerability analysis
      • Possible to repair in place: Estimate: 1 week
      • Unable to repair in place. Column probably needs to be removed from the wall: Estimate: 1 month
  • Low Energy
    • 750 Buncher gradient bouncing problem
      • Replaced IPA2 tube
      • Station ran for 1 hr and failed hard
      • Extensive troubleshooting on the station followed
      • Found a 1 mH inductor blown up on the 7651 control grid socket
      • After replacement, the station has been running without problems
  • High Energy
    • Replaced a HE trim steering magnet chassis/power supply
    • Installed new fans on all the 500W Amplifier modules
    • Water group replaced a defected turbine on skid#3 for the chilled water
    • Downloaded new VXI software on all HE Klystrons
  • Next Week
    • Support Neutron Therapy Facility patient treatment (Tuesday, Thursday, Friday)
    • Address problems as they reveal.
    • Provide beam to downstream machines.
    • Continue machine tuning

Booster

Shutdown Week 10 – On Schedule

  • Tunnel Activities
    • Running beam 1-3 turns, 85% efficiency
    • Systematic checkout of all the new correctors, BPMs, and BLMs
    • Accessed tunnel Wednesday for BRF9 PA replacement
  • Next Week’s plan
    • Continue tuning, reducing losses, improving efficiencies, and increasing intensity

Antiproton Source

Transition from Shutdown to Start-up

  • Terminated bake on Debuncher injection septum Saturday
    • X20 improvement in vacuum
  • Tunnels secured late Sunday and Monday
  • Safety System testing on Monday
  • Power supply testing since Tuesday
    • Problems with A:IB, D:IB, A:LQ and D:QF fixed
  • Occasional accesses this week

Startup Plan

  • When 8-GeV protons are available from MI
  • When 120-GeV protons are available from MI
    • Target station check-out
    • Beam to Debuncher
      • Check out RF
      • First pass at Debuncher cooling phasing (when cold)
    • Beam to Accumulator, stack pbars
      • Phase 4-8 GHz core system
    • Transfer antiprotons to Recycler
  • When stable operation resumes
    • Full check-out of all cooling systems

Main Injector

Tunnel Work

MI-8

  • Bake pulled on Wednesday
  • MI-8 line work complete

MI-30

  • Vacuum leak repaired and ion pumps replaced
  • Masks installed, sand removed but replacement shielding at the primary in progress

MI-40

  • Kicker vacuum leak check complete. One new 150L ion pump failed and replaced requiring a re-bake of the RGA manifold
  • Lambertson bake finished. There is a vacuum leak on the large flange

MI eCloud (MI-52)

  • Work complete

MI RF

  • Repairs to South Anode PS appear successful
  • Vacuum issues @604 being addressed

MIPS work wrapping up.

LCW systems on and ready

New Fluorinert skid installed for MI-304 kicker.

Today

  • Finish up MI-40 Kickers
  • Deal with our newest vacuum leak (Lam40)
  • Finish primary collimator shielding work
  • Clean up tunnel
  • Final walk throughs
  • Clear bus faults
  • Install sand shielding @ MI-14 and MI-39

Tonight

  • System checks

Saturday/Sunday

  • Access will be needed to repair Lambertson vacuum leak
  • Hipot
  • Energize ramp

Recycler

Recycler

  • RR129 & RR400 repair/modifications ( Complete )
  • HLS installation at MI-10 ( Complete)
    • Last miscellaneous tweaks
  • Vacuum maintenance
    • Need to sublimate TSPs and degas ion pumps
      • Critical region is 103-109
    • One more assessment on Monday with MI ramping
      • Take advantage of the (very) probable access mid-week for alignment
  • Alignment finishing up (?)
  • 1-2 GHz stochastic cooling system converted back to longitudinal
  • CIA crate activities ( Complete)
  • Power supplies maintenance ( Complete)
  • Move 6 magnets ( complete)
  • RF back on
    • Tested okay
    • RF states cleanup underway
  • Correctors PS on
    • Ramps loaded and checked
  • Kickers tested
    • One small timing problem was fixed
  • Walk-through today
  • Check movable devices today
    • Scrapers (including ECool's)
    • Stochastic pickups & kickers
  • Check RF & LLRF files
    • In progress

Electron Cooling

  • Failed components inside the Pelletron fixed on Sunday
  • Transfer gas on Monday
  • Problems with the new NESLAB unit (heat exchanger)
    • Taken out of the system
    • Test on the bench
  • Conditioned HV section-by-section
  • 5.2 MV all sections together
  • Installed new 2” GVM and checked calibration
    • Off by ~2.6%
  • Established recirculating beam
    • 200 mA
  • Pelletron maintenance ( Complete)
  • Water/Glycol system cleanup and upgrade ( Complete)
  • NMR probes replacement ( Complete)

Tevatron

Other Items

  • D-1 passed Hipot locks on cryo box removed today
  • Finishing details in E-1 ready for Hipot sign off Tuesday
  • E0 vacuum work done
  • B3 vacuum OK
  • B 27 house needs new turbo
  • A4 CV3 U&D have water in them work in progress.
  • Vertical crystal collimator controls cables installed termination in progress
  • Poly bags installed at B0 and D0 radiation safety can take control of D0
  • TeV re-cored to running keys
  • Start TeV Sequencer Monday

To Do List

  • Finished B0 HLS
  • Install luminosity counters at B0
  • Change locks on panels at B0 to Rad safety locks
  • Remove golf cart

NuMI

Shutdown Week 11

  • Test Beam Loss Monitors
  • Check Out Baffle Thermocouples
  • Completed Horn Pulse Testing & Thermal Imaging of Stripline in Anti-Neutrino Mode
  • Install & Seal Target Pile R-Blocks Over Horns 1&2
  • Dry set R-Blocks over Target
  • Re-Cored Target Hall & Absorber Hall
  • Convert Horn Pulse Power Supply back to Neutrino Mode
  • Replace Desiccant Wheel Drive Belts on Target Pile Dehumidifiers
  • Complete Horn Pulse Testing in Neutrino Mode
  • EPB Dipole PEEK Installation Complete
  • Safety Interlock Testing Complete
  • Target & Absorber Hall Clean Up Complete
  • Installed T-Hall Shielding Plug & Air Barrier
  • Water Sample Taken from Drainage System Pre Target to Absorber Hall (16 samples pts.)
  • Commission Absorber Hall Dehumidification System
  • Replace Damaged Temperature Controllers in Absorber Hall Dehumidifiers
  • Take Decay Pipe Helium Gas Samples
  • Install New Decay Pipe Pressure Relief Valves
  • Possible leak on reservoir tank located on vacuum skid
  • Repair Control Valve on Absorber Intermediate Skid
  • Clean out Pre-Target, Target Hall, & Absorber Hall

Low Intensity Run Plan

Day One

  • 1300 Target Scan
  • 1600 Realign Target Transversely
  • 2000 Target Scan
  • 2300 Horn Current Scan
  • 0100 Put Absorber in Front of Muon Monitor
  • 0200 Horn Current Scan
  • 0400 Day Over

Day Two

  • 0700 Move Target Longitudinally
  • 1300 Target Scan
  • 1600 Realign Target Transversely
  • 2000 Target Scan
  • 2300 Horn Current Scan
  • 0100 Put Absorber in Front of Muon Monitor
  • 0200 Horn Current Scan
  • 0400 Day Over

Day Three

  • 0700 Move Target Longitudinally
  • 1300 Target Scan
  • 1600 Realign Target Transversely
  • 2000 Target Scan
  • 2300 Horn Current Scan
  • 0100 Put Absorber in Front of Muon Monitor
  • 0200 Horn Current Scan
  • 0400 Day Over

Day Four

  • 0700 Move Target Longitudinally
  • 1300 Target Scan
  • 1600 Realign Target Transversely
  • 2000 Target Scan
  • 2300 Horn Current Scan
  • 0100 Put Absorber in Front of Muon Monitor
  • 0200 Horn Current Scan
  • 0400 Run Over

Day Five (Target Hall Access)

  • Seal R-Blocks over Target on Target Pile
  • Let Caulk Dry Over Night

Day 6 (Target Hall Access)

  • Remove Crane Electronics
  • Ramp up Target Pile Fan
  • Remove Crane Electronics

Cryogenic System

TeV at LN2 Temperatures

Tevatron Operations and Maintenance Status

  • Continued support AD/MSD for vacuum repairs: D3, E3, F1, A3, A2, and Pbar Accumulator tanks are back at LN2 temperature. A1 was being cooled down to LN2 temperature but was halted due to contamination plugging of heat exchanger in valve box. D1 and E1 under LOTO for repairs by MSD.
  • LN2 inventory control and controlling magnet temperatures at LN2.
  • Contamination plugging of heat exchangers occurred starting Monday morning; first at A1 in valve box, then F3 in HTX-1, F1 in HTX-1, and partial plugging in E4 HTX-1. These were derimed per procedure used at end of 2007 shutdown period when similar plugging occurred at A3, F3, and F4. Contamination source is suspected as oil volatiles from A0#1 Mycom compressor; it has been isolated and further investigation/resolution continues.
  • Ongoing numerous FRIG/Compressor controls/instrumentation maintenance/repairs.
  • Vacuum pumping of various cryo equipment in FRIG buildings to start next week

---------------------------------------------------------------------------------------

Tevatron cryo schedule is dependent on when we start system purification, which is awaiting resolution of htx plugging issue.

CHL HELIUM and NITROGEN PLANTS

  • 99.9% of all scheduled routine mechanical and electrical maintenance is complete

Current operations:

  • Compressor “B” and “D” were purified, as well as Helium Coldbox-2
  • Helium Coldbox-2 was cooled down and in process of cooling 10,000-gallon LHe storage tank.
  • Bulk LHe delivery is tentatively scheduled for Monday, in anticipation of Tevatron FRIG system cooldown subsequently.
  • Mobile Helium Purifier at CHL has been set up for Tevatron FRIG purification, awaiting start of purification

-----------------------------------------------------------------------------------------------------

We're back in full 24/7 shift operations, supporting Tevatron FRIG system monitoring and operations, as well as A0 Test Caves and such.

Controls

  • New Linux Kernels

MiniBooNE

  • All systems are up and running except for the critical devices and magnets

Collider Detector at Fermilab

  • Muon high voltage and silicon power supply work continues

EE Support

  • Finishing up MI power supply work
  • Starting TeV work next week

MTA

  • The permanent magnet has been installed, but needs to be aligned

FESS

Electrical and High Voltage

  • Kautz Rd SS – work completed & on line.
  • MR – 30 hp pond pump/motors PMs scheduled for Sat 8/29.
  • TeV Lighting CWO scheduled
  • MI40 exhaust fan motor to be replaced
  • MI 20 exhaust fan wiring/conduit repair scheduled.
  • MI Lighting CWO scheduled.
  • Switchyard - new air switch installed

Mechanical

  • Booster ring tunnel – sump pump valve & fitting leak repair to be scheduled. Exhaust fan needs belts requiring access/scaffold.
  • CDF/BO – supply air handlers[AC-9] – sheave belt alignment & coil cleaning.
  • AP 20 & 40 – exhaust fan sheaves replacement to be scheduled.
  • MR – F-3 [30 hp] pond bypass pump & check valve replacement.
  • HVAC
  • MI – AC Marvair checks scheduled for next week

Chiller #3 Update

  • They will have all the tubes[~1000] through the first head and supports by end of day today
  • Will weld on the 2nd head tomorrow & then start expanding the tubes in place.
  • They hope to beat the 9/14 date but will depend on the leak check & refrigerant charging & final commissioning

ICW

  • Spot flushing – done on 8/25 with many ZMs flushed out on Meson line – will need to schedule another spot flush. Strainer checks were ok

CUB

  • 55LCW chiller # 3 – tube & header work progressing – all old tubes/headers removed & new tubes being installed. Work on schedule to be completed by 9/14 or sooner.
  • Comfort chiller # 5 – refrigerant & prep work taking longer than expected. Should be ready by mid next week

Switchyard

D-Zero

  • Their detector has been closed

Mechanical

  • Moving out of MI
  • The MI Lambertson has a vacuum leak

Instrumentation

  • Loss monitor work in MI
  • BPM work in the Recycler

Alignment

Alignment

  • Stickmic survey of Muon system including CF and EFs after closing detector. The start and finish dates are preliminary and depend on shutdown progress.
  • Align 2 EPB dipoles in enclosure B (VH94) before reworked EPB dipoles. Align 2 new dipoles
  • Align new permanent magnet to be installed on the Linac side
  • Replace/ Install SEM AP0

Safety

Safety

  • Accelerator restart sheets are available from Mike Gerardi

Interlocks

  • Misc. Tevatron Interlock repairs & ground fault diagnostics installation complete
  • MINOS Absorber interlock modifications complete
  • Safety System Testing
    • Interlocks tested this week:
      • MI, NuMI & F-Sector Enclosures & Radiation Monitors, Recycler Coasting Beam System, Pelletron and Booster
      • F-Sector ODH Monitors
    • Interlock testing next week:
      • CDC testing for MI, P-Bar, NuMI, and MiniBooNE this weekend
      • Collision Halls, Transfer Hall, A-E Sector Enclosures
      • TH & A-E Sector ODH Monitors

Continue using the Take Five principles before, during, and after work activities to:

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 08/28/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory